I need to come up some equations to this problem, and Im having a little trouble.
A sales representative informs you that there are two floor plans still available and that there are a total of 56 houses available. Use x to represent floor plan #1 and y to represent floor plan #2 Write an equations that illustrates the situation.
Then write another equation indicating that there are 3 times as many home available with the second floor plan than the first, using the same variables.

Answer by Theo(13342)

x = number of floor plan 1
y = number of floor plan 2

total number of houses = 56 with 1 floor plan per house.

x + y = 56

number of floor plan 2 = 3 * number of floor plan 1

y = number of floor plan 2 so y = 3x

your equations are:

x + y = 56
y = 3x

replace y with 3x in first equation to get:

x + 3x = 56
this becomes:
4x = 56 which becomes:
x = 14
this makes y = 3x = 42

42 + 14 = 56 so number of total floor plans is correct.
x = number of floor plan 1 = 14
y = number of floor plan 2 = 42